6. (Daten precedes thorms on Jupiter have noticed that a low pressure (Data on front page may help you.) (4 points) NASA scientists studying Jup x 10 torr storms. The atmospheric pressure on Jupiter is 5.57 a. Calculate the pressure in atm. b. Calculate the pressure in psi. 7(6 points) if a robotic vehicle were sent to Jupiter and the tires were filled with krypton, answer the following questions about the tires Jupiter where the temperature is -108°C, what would the volume of krypton in the tires be? a. If the tires were filled with 2000. g of krypton at 842 atm, a typical pressure on During a storm, the tire pressure decreases from 842 atm to 647 atm and the temperature increases from -108°C to -84°C. If the original volume of a different set of tires was 2.18 L, what would the volume of the tires be during the storm? b.
Answer #1

6) a) We calculate the pressures to different units:

P = 5.57x10 ^ 5 Torr * (1 atm / 760 Torr) = 732.89 atm

b) P = 5.57x10 ^ 5 Torr * (1 psi / 51.71 Torr) = 10771.61 psi

7) a) Calculate the moles of Kripton:

n Kr = 2000 gr * (1 mol / 83.8 g) = 23.87 moles

T = -108 ºC + 273 = 165 K

Using the ideal gas equation, calculate the volume:

V = n * R * T / P = 23.87 * 0.082 * 165/842 = 0.384 L

b) by ideal gas law we have:

P1 * V1 * T2 = P2 * V2 * T1

Clearing final volume:

V2 = P1 * V1 * T2 / P2 * T1 = 842 * 2.18 * 189/647 * 165 = 3.25 L